OpenStudy (anonymous):

A high-altitude spherical weather balloon expands as it rises due to the drop in atmospheric pressure. Suppose that the radius r increases at the rate of 0.16 inches per second and that r = 38 inches at time t = 0. Determine the equation that models the volume V of the balloon at time t and find the volume when t = 280 seconds.

OpenStudy (nincompoop):

you also have to understand that the temperature is another variable to consider.

OpenStudy (anonymous):

r(t)=0.16*t+38 V=(4/3)(pi)(r(t))^3 That is where I am going...

OpenStudy (anonymous):

V(t) = (0.16t +38)^3 * 4/3 * pi

OpenStudy (anonymous):

Volume as a function of time: V(t)

OpenStudy (anonymous):

This is my final answer: 

OpenStudy (anonymous):

What is 2,377,823,51 in^3?

OpenStudy (anonymous):

t = 280 yes, your answer is correct

OpenStudy (anonymous):

that is a huge balloon

OpenStudy (anonymous):

I know, laughing out loud ! Thank for your help :)

OpenStudy (anonymous):

I meant thank You

OpenStudy (anonymous):

No problem